find the inverse function g of the given function and state the domain and range of g
y=f(x)=(x+2)^3

Answer by edjones(8007)

y=f(x)=(x+2)^3
x=(y+2)^3 switch the x and y and solve for y.
x^1/3=y+2 take cube root of both sides.
y=x^1/3 -2
f(g)=x^1/3 -2
The domain and range are all real numbers.
Ed